OpenStudy (moongazer):

If y^3 = at^2 , then (d^2 y)/dt^2 ?

OpenStudy (moongazer):

I only got to the part: dy/dt = 2at/3y^2

OpenStudy (yttrium):

Just derive again, so you can solve the d^2y/dt^2. :))

OpenStudy (yttrium):

it's like solve for y''

OpenStudy (amriju):

let it remain as 3y^2=2at..differntiate again....using multiplication rule....

OpenStudy (moongazer):

it says the answer is: -2a/9y^2 but I am getting a different answer

OpenStudy (moongazer):

is "a" here constant?

OpenStudy (yttrium):

Quotient rule, remember?

OpenStudy (amriju):

yo constnt

OpenStudy (yttrium):

for easier calculation, you can factor out 2a/3 and just derive (t/y^2)

OpenStudy (moongazer):

could you show me the first step for the second derivative? maybe I did something wrong

OpenStudy (moongazer):

i'll try again :)

OpenStudy (amriju):

may be u need to substitute y as (at^2)^2/3

OpenStudy (moongazer):

I got: (6ay^4 - 8a^2 t^2 y)/9y^6 I think I am doing something wrong.

OpenStudy (moongazer):

i'll try again :)

OpenStudy (yttrium):

\[y' = \frac{ 2at }{ 3y^2 }\], right? Therefore, \[y' = \frac{ 2a }{ 3 } (\frac{ t }{ y^2 })\] \[y'' = \frac{ 2a }{ 3 } [\frac{ y^2 - 2tyy') }{ y^4 }]\] \[y'' = \frac{ 2a }{ 3 } \frac{ y^2 - 2ty(\frac{ 2a }{ 3 })(\frac{ t }{ y^2 }) }{ y^4 }\] Simplifying this equation we will get. \[y'' = \frac{ 2a }{ 3 } [ \frac{ 3y^4 -4at^2y }{ 3y^6 }]\]

OpenStudy (yttrium):

can you follow @moongazer ?

OpenStudy (moongazer):

That's what I got. I think there is some typo in my answer sheet. Thanks :)

myininaya (myininaya):

Well you could simplify

OpenStudy (yttrium):

@moongazer, it isn't the answer in simplest form. Maybe you can continue this and let us verify if we get the same answers.

OpenStudy (moongazer):

@myininaya could you simply it further? @Yttrium yes, I understood your solution

myininaya (myininaya):

Could I? Yes. Can you?

myininaya (myininaya):

Eye ball the numerator and the denominator. You should see they share a common factor.

OpenStudy (yttrium):

For our class, the definition of simplest form is that it is factored out completely. Just factor out some stuff and do cancellation.

OpenStudy (moongazer):

2a(3y^3 - 4at^2) / 9y^5

OpenStudy (moongazer):

that's what I got

myininaya (myininaya):

You can also choose to write at^2 as y^3

myininaya (myininaya):

Recall our initial equation: y^3=at^2

myininaya (myininaya):

I see at^2

myininaya (myininaya):

Replace it y^3 then combine like terms

OpenStudy (yttrium):

@moongazer , do you what myininaya told you?

myininaya (myininaya):

@moongazer I'm leaving for tonight but I think Yttrium is still here to help. Goodnight you guys. I think you guys got it from here for sure. :)

OpenStudy (moongazer):

WOW! Thank you very much to both of you. I failed to notice that until you said it. Thanks :) I now got the answer. :)